# BeeManc at the PLABA Track of TAC-2024: RoBERTa for Task 1 – LLaMA3.1 and GPT-4o for Task 2

### v2

@@ Abstract
This report is the system description of the BeeManc team for shared task Plain Language Adaptation of Biomedical Abstracts (PLABA) 2024. This report contains two sections corresponding to the two sub-tasks in PLABA-2024. In task one, we applied fine-tuned ReBERTa-Base models to identify and classify the difficult terms, jargon and acronyms in the biomedical abstracts and reported the F1 score. Due to time constraints, we didn’t finish the replacement task. In task two, we leveraged Llamma3.1-70B-Instruct and GPT-4o with the one-shot prompts to complete the abstract adaptation and reported the scores in BLEU, SARI, BERTScore, LENS, and SALSA. From the official Evaluation from PLABA-2024 on Task 1A and 1B, our much smaller fine-tuned RoBERTa-Base model ranked 3rd and 2nd respectively on the two sub-tasks, and the 1st on averaged F1 scores across the two tasks from 9 evaluated systems. Our LLaMA-3.1-70B-instructed model achieved the highest Completeness score for Task-2. We share our source codes, fine-tuned models, and related resources at https://github.com/HECTA-UoM/PLABA-2024
@@ 5 Conclusion and Future Work
In this report, we present the two systems submitted by the BeeManc team for two different tasks in PLABA-2024. For Task 1, we submitted a classification system based on Roberta for Task 1a and 1b. The system’s performance at the word level was not outstanding, highlighting the limitations of simple classification models in multi-label, multi-class tasks. For Task 2, we submitted end-to-end generation results based on LLaMa 3.1 and GPT-4o for the generation task. On the Evaluation set, GPT-4o performed slightly better than LLaMa. In the unsupervised evaluation on the test set, GPT-4o also outperformed LLaMa 3.1. Our findings demonstrate the potential of open-source models in text simplification tasks within the medical domain.
@@ 6 Acknowledgment
We thank Nicolo and Sam for their valuable discussion and feedback during this project with their expertise from the experience of PLABA-2023. The authors wrote the initial draft, used Generative AI to revise/polish it, then the authors reviewed the generated content. The authors used ChatGPT to refine the paper writing.

### v4

# MaLei at the PLABA Track of TREC 2024: RoBERTa for Term Replacement – LLaMA3.1 and GPT-4o for Complete Abstract Adaptation

@@ Abstract
This report is the system description of the MaLei team (Manchester and Leiden) for the shared task Plain Language Adaptation of Biomedical Abstracts (PLABA) 2024 (we had an earlier name BeeManc following last year), affiliated with TREC2024 (33rd Text REtrieval Conference https://ir.nist.gov/evalbase/conf/trec-2024). This report contains two sections corresponding to the two sub-tasks in PLABA-2024. In task one (term replacement), we applied fine-tuned ReBERTa-Base models to identify and classify the difficult terms, jargon, and acronyms in the biomedical abstracts and reported the F1 score (Task 1A and 1B). In task two (complete abstract adaptation), we leveraged Llamma3.1-70B-Instruct and GPT-4o with the one-shot prompts to complete the abstract adaptation and reported the scores in BLEU, SARI, BERTScore, LENS, and SALSA. From the official Evaluation from PLABA-2024 on Task 1A and 1B, our much smaller fine-tuned RoBERTa-Base model ranked 3rd and 2nd respectively on the two sub-tasks, and the 1st on averaged F1 scores across the two tasks from 9 evaluated systems. Our LLaMA-3.1-70B-instructed model achieved the highest Completeness score for Task 2. We share our source codes, fine-tuned models, and related resources at https://github.com/HECTA-UoM/PLABA2024
@@ 5 Conclusion and Future Work
In this report, we present the two systems submitted by the MaLei team for two different tasks in PLABA-2024. For Task 1, we submitted a classification system based on Roberta for Task 1a and 1b. The system’s performance at the word level was not outstanding, highlighting the limitations of simple classification models in multi-label, multi-class tasks. For Task 2, we submitted end-to-end generation results based on LLaMa 3.1 and GPT-4o for the generation task. On the Evaluation set, GPT-4o performed slightly better than LLaMa. In the unsupervised evaluation on the test set, GPT-4o also outperformed LLaMa 3.1. Our findings demonstrate the potential of open-source models in text simplification tasks within the medical domain.
@@ 6 Acknowledgment
We thank Nicolo and Sam for their valuable discussion and feedback during this project with their expertise from the experience of PLABA-2023. The authors wrote the initial draft, used Generative AI to revise/polish it, then the authors reviewed the generated content. The authors used ChatGPT to refine the paper writing. LH and GN were supported by the grant “Integrating hospital outpatient letters into the healthcare data space” (EP/V047949/1; funder: UKRI/EPSRC). LH is also grateful to the EU grant 4D Pictures https://4dpicture.eu/.